# Some Primal-Dual Theory for Subgradient Methods for Strongly Convex Optimization

### v3

@@ 2.1 Related Work
Switching Subgradient Method Guarantees Convergence rate guarantees for the switching subgradient method of [10] have been extensively studied for convex Lipschitz minimization [19, 20, 21, 22] and more recently for non-Lipschitz settings [23]. Our theory extends this prior theory to give matching dual bounds and identifies Lagrange multipliers $u_{s}=\frac{\sum_{k:s(y_{k})=s}\lambda_{k}}{\sum_{k:s(y_{k})=0}\lambda_{k}}$ implicitly built by this classic method (at no added cost). Recently, nonconvex guarantees were developed by [24] but are beyond our scope.
@@ 2.2 Assumptions for our Convergence Theory
Assumption B. There exists some $x_{\mathtt{SL}}\in\mathrm{dom\ }\partial r$ with $f_{s}(x)<0$ for all $s=1\dots m$.
Note strong convexity ensures there exists a unique minimizer $x_{\mathtt{OPT}}\in\mathrm{dom\ }\partial r$ of (1.1). These two points $x_{\mathtt{OPT}},x_{\mathtt{SL}}\in\mathrm{dom\ }\partial r$ serve as important references that our analysis is done with respect to. We fix two subgradients of $r$ at these points: The subgradient $n_{x_{\mathtt{OPT}}}\in\partial r(x_{\mathtt{OPT}})$ is chosen such that $f(x)+r(x_{\mathtt{OPT}})+\langle n_{x_{\mathtt{OPT}}},x-x_{\mathtt{OPT}}\rangle$ minimizes over $f_{s}(x)\leq 0$ at $x_{\mathtt{OPT}}$. The subgradient $n_{x_{\mathtt{SL}}}\in\partial r(x_{\mathtt{SL}})$ can be chosen freely.
These two reference subgradients of $r$ facilitate considering two lower bounding relaxations of the objective $f_{0}+r$ for either $y\in\{x_{\mathtt{OPT}},x_{\mathtt{SL}}\}$, denoted by
@@
We denote the (relaxed) objective gap with respect to a reference point $y\in\{x_{\mathtt{OPT}},x_{\mathtt{SL}}\}$ at each iteration on the selected function $f_{s(x_{k})}$ by
@@
To see the necessity of using a relaxation, consider $r$ as an indicator function for some simple constraint set. This set is projected onto each iteration where $x_{k}$ is feasible for all of the functional constraints, ensuring $x_{k+1}$ is feasible for the simple constraint. As a result, we cannot guarantee that $x_{k}$ satisfies the simple constraint or that $x_{k+1}$ satisfies the functional constraints. So, some notion of feasibility must be relaxed.
@@
Assumption C. For both $y\in\{x_{\mathtt{OPT}},x_{\mathtt{SL}}\}$, every iterate $x_{k}$ has
@@
If one sets $L_{1}=0$, this amounts to assuming a uniform bound on subgradients. As a more interesting example, this model captures the minimization of any strongly convex sum $h=h_{1}+h_{2}$ where $h_{1}$ is $L$-smooth and $h_{2}$ is nonsmooth but $M$-Lipschitz. In this case, although $h$ may be non-Lipschitz, every subgradient $g\in\partial h(x)$ has [^3] As a short derivation of this, by the sum rule of subgradient calculus, let $g_{i}\in\partial h_{i}(x)$ have $g=g_{1}+g_{2}$. Since $h$ is strongly convex, let $x^{*}$ denote its unique minimizer with subgradients $g^{*}_{i}\in\partial h_{i}(x^{*})$ satisfying $g_{1}^{*}+g_{2}^{*}=0$. Then $\|g\|_{2}^{2}\leq 2\|g_{2}-g_{2}^{*}\|_{2}^{2}+2\|g_{1}+g^{*}_{2}\|_{2}^{2}\leq 8M+4L(h(x)-h(x^{*}))$ where the first inequality bounds $\|a+b\|_{2}^{2}\leq 2\|a\|^{2}_{2}+2\|b\|^{2}_{2}$ and the second applies the Lipschitz continuity of $h_{2}$ to bound $\|g_{2}-g^{*}_{2}\|_{2}\leq 2M$ and the smoothness of $h_{1}+\langle g^{*}_{2},x\rangle$ to bound $\|g_{1}+g^{*}_{2}\|_{2}^{2}\leq 2L(h_{1}(x)-h_{1}(x^{*})+\langle g_{2}^{*},x-x^{*}\rangle)\leq 2L(h(x)-h(x^{*}))$.
$\|g\|_{2}^{2}\leq 8M^{2}+4L(h(x)-\inf h)\ .$ (2.9)
Section 4 gives an illustrative numerical example of this form.
@@
$T_{0}$ $:=\sup\left\{k\in\mathbb{N}\mid\alpha_{k}>1/L_{1}\right\}\ ,$
$C_{0}$ $:=\sum_{k=0}^{T_{0}}\lambda_{k}\max\left\{L_{1}\alpha_{k}-1,0\right\}\max\{\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}}),\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{SL}})\}\ .$
Observe that $T_{0}$, and hence $C_{0}$, is bounded if $\alpha_{k}$ is eventually always less than $1/L_{1}$, capturing all stepsize policies with $\alpha_{k}\rightarrow 0$. Despite being bounded, the constant $C_{0}$ can be exponentially large in $T_{0}$. The toy example considered above has $T_{0}=397$ and $C_{0}>10^{112}$. Such potentially exponential-sized constants can be avoided through careful stepsize selection as if one selects $\alpha_{k}\leq 1/L_{1}$, $T_{0}=-\infty$ and $C_{0}=0$. Regardless, our convergence guarantees apply whenever $C_{0}$ is finite. As a natural consequence of our main convergence analysis, we find the rate $x_{k}$ can diverge, and hence the constant $C_{0}$, are at most exponential in $T_{0}$ (see Proposition 3.4).
@@ 3 Primal-Dual Equivalence and Convergence Analysis
Let $x_{k}$ be the sequence of iterates of the primal method (2.2) with stepsizes $\alpha_{k}$ and $y_{k}$ be the sequence of iterates of the dual method (2.3) for some $\lambda_{k},\beta_{k},\mu\geq 0$. If $x_{0}=y_{0}$, $\beta_{k}=\bar{\beta}\geq 0$ is constant, and $\alpha_{k}=\lambda_{k}/(\mu\sum_{i=0}^{k}\lambda_{i}+\bar{\beta})$, then these methods are equivalent, that is $x_{k}=y_{k}$.
@@
By assumption, $x_{0}=y_{0}$. Supposing $x_{i}=y_{i}$ for all $i=0,\dots,k$, we will inductively prove the result by showing $x_{k+1}=y_{k+1}$. Let $\bar{g}_{i}=g_{0}(y_{i};\xi_{i})+n_{i+1}$ if $y_{i}$ is feasible and $g_{s(y_{i})}(y_{i};\xi_{i})$ otherwise. For each $i<k$, we claim $x_{i+1}=x_{i}-\alpha_{i}\bar{g}_{i}$. If $x_{i}$ is infeasible, this is immediate. If $x_{i}$ is feasible, since $x_{i+1}=y_{i+1}$, we have $n_{i+1}\in\partial r(x_{i+1})$. As a result, $x_{i+1}=x_{i}-\alpha_{i}\bar{g}_{i}$ by (2.1). Inductively, we conclude for $i<k$ that $x_{i+1}-x_{0}=-\sum_{t=0}^{i}\alpha_{t}\bar{g}_{t}.$
By Lemma 2.1, $y_{k+1}$ is the unique solution to $\sum_{i=0}^{k}\lambda_{i}\left(\bar{g}_{i}+\mu(y_{k+1}-y_{i})\right)+\beta(y_{k+1}-y_{0})=0.$ Rearranging and simplifying this, it follows that
$y_{k+1}$ $=y_{0}-\frac{\sum_{i=0}^{k}\lambda_{i}\mu(y_{i}-y_{0})-\sum_{i=0}^{k}\lambda_{i}\bar{g}_{i}}{\sum_{i=0}^{k}\lambda_{i}\mu+\bar{\beta}}$
@@
$=y_{0}-\frac{\sum_{t=0}^{k-1}\lambda_{t}\bar{g}_{t}\frac{\sum_{i=t+1}^{k}\mu\lambda_{i}}{\sum_{i=0}^{t}\mu\lambda_{t}+\bar{\beta}}+\sum_{i=0}^{k}\lambda_{i}\bar{g}_{i}}{\sum_{i=0}^{k}\lambda_{i}\mu+\bar{\beta}}$
$=y_{0}-\frac{\sum_{t=0}^{k-1}\lambda_{t}\bar{g}_{t}\left(\frac{\sum_{i=t+1}^{k}\mu\lambda_{i}}{\sum_{i=0}^{t}\mu\lambda_{t}+\bar{\beta}}+1\right)}{\sum_{i=0}^{k}\lambda_{i}\mu+\bar{\beta}}-\frac{\lambda_{k}\bar{g}_{k}}{\sum_{i=0}^{k}\lambda_{i}\mu+\bar{\beta}}$
$=y_{0}-\sum_{t=0}^{k-1}\frac{\lambda_{t}\bar{g}_{t}}{\sum_{i=0}^{t}\lambda_{t}\mu+\bar{\beta}}-\frac{\lambda_{k}\bar{g}_{k}}{\sum_{i=0}^{k}\lambda_{i}\mu+\bar{\beta}}$
@@
where the second equality uses the inductive hypothesis for each $y_{i}-y_{0}=x_{i}-x_{0}=-\sum_{t=0}^{i-1}\alpha_{t}\bar{g}_{t}$, the third exchanges summands, and the remainder combine and simplify terms. If $x_{k}$ is infeasible, its immediate that $x_{k+1}=x_{k}-\alpha_{k}\bar{g}_{k}=y_{k+1}$. If $x_{k}$ is feasible, we know $n_{k+1}\in\partial r(x_{k}-\alpha_{k}\bar{g}_{k})$. Then by (2.1), $x_{k+1}=x_{k}-\alpha_{k}\bar{g}_{k}=y_{k+1}$. ∎
@@
Hence, provided dual weights $\lambda_{k}$, one can easily construct $\alpha_{k}$ as stated in the theorem, and conversely, given stepsizes $\alpha_{k}$, one can easily construct corresponding weights $\lambda_{k}$. For example, fixing $\lambda_{0}=1$, multipliers for several common stepsizes are given below
@@ 3.1 Statement of Primal-Dual Convergence Guarantees
$\mathtt{primal\mbox{-}gap}_{T}:=\frac{\sum_{k<T:s(x_{k})=0}\lambda_{k}h_{x_{\mathtt{OPT}}}(x_{k})}{\sum_{k<T:s(x_{k})=0}\lambda_{k}}-p_{\star}\ ,$
@@
$\mathtt{dual\mbox{-}gap}_{T}:=p_{\star}-\inf\frac{M^{(T-1)}}{\sum_{k<T:s(x_{k})=0}\lambda_{k}}\ ,$
@@
The stochastic switching proximal subgradient method (2.2) has
$\mathbb{E}_{\xi}\left[\frac{\sum_{k<T:s(x_{k})=0}\lambda_{k}}{\sum_{k=0}^{T-1}\lambda_{k}}\right]\geq\frac{\tau_{\mathtt{SL}}}{2(h_{x_{\mathtt{SL}}}(x_{\mathtt{SL}})-\inf h_{x_{\mathtt{SL}}})+\tau_{\mathtt{SL}}}\left(1-\frac{L_{0}^{2}\sum_{k=0}^{T-1}\lambda_{k}\alpha_{k}+C_{0}}{\sum_{k=0}^{T-1}\lambda_{k}}\right)$
@@
Under Assumptions A-C, the stochastic switching proximal subgradient method (2.2) with stepsizes $\alpha_{k}>0$ and associated dual weights $\lambda_{k}>0$ and $\beta_{k}=0$ has
@@
Theorem 3.2 recovers the primal convergence rate of [11, Section 3.2] with $\alpha_{k}=2/(\mu(k+2))$ and extends it to be a primal-dual guarantee covering proximal, switching, and non-Lipschitz settings. Theorem 3.1 shows this stepsize corresponds to dual averaging with weights $\lambda_{k}=k+1$. When $m=0$, Theorem 3.2 ensures
@@
using that $\sum_{k=0}^{T-1}\lambda_{k}=T(T+1)/2$ and $\sum_{k=0}^{T-1}\lambda_{k}\alpha_{k}=2(T-\sum_{k=2}^{T}1/k)/\mu\leq 2T/\mu$. When $m>0$ and the subgradient oracle is deterministic, applying Proposition 3.1 gives a rate worse by only a factor depending on the Slater point of
$\frac{2(h_{x_{\mathtt{SL}}}(x_{\mathtt{SL}})-\inf h_{x_{\mathtt{SL}}})+\tau_{\mathtt{SL}}}{\tau_{\mathtt{SL}}\left(1-\frac{4L_{0}^{2}}{\mu(T+1)}-\frac{2C_{0}}{T(T+1)}\right)}\left(\frac{4L_{0}^{2}}{\mu(T+1)}+\frac{2C_{0}}{T(T+1)}\right)\ .$
Here, the role of $C_{0}$ bounding the effect of the non-Lipschitz constant $L_{1}$ becomes clear. The only role $L_{1}$ plays in our rate via $T_{0}$, which in turn $C_{0}$ may be exponentially large in (see Proposition 3.4). If $C_{0}$ is small, then convergence will be dominated by the classic $O(L_{0}^{2}/\mu T)$ term as the dependence on $C_{0}$ shrinks at a fast $O(1/T^{2})$ rate.

### v5

@@ 2.1 Related Work
Switching Subgradient Method Guarantees. Convergence rate guarantees for the switching subgradient method of [10] have been extensively studied for convex Lipschitz minimization [19, 20, 21, 22] and more recently for non-Lipschitz settings [23]. Our theory extends this prior theory to give matching dual bounds and identifies Lagrange multipliers $u_{s}=\frac{\sum_{k:s(y_{k})=s}\lambda_{k}}{\sum_{k:s(y_{k})=0}\lambda_{k}}$ implicitly built by this classic method (at no added cost). Recently, nonconvex guarantees were developed by [24] but are beyond our scope.
@@ 2.2 Assumptions for our Convergence Theory
Assumption B. There exists some $x_{\mathtt{SL}}\in\mathrm{dom\ }\partial r$ with $f_{s}(x_{\mathtt{SL}})<0$ for all $s=1\dots m$.
Note strong convexity ensures there exists a unique minimizer $x_{\mathtt{OPT}}\in\mathrm{dom\ }\partial r$ of (1.1). These two points $x_{\mathtt{OPT}},x_{\mathtt{SL}}\in\mathrm{dom\ }\partial r$ serve as important references that our analysis is done with respect to. We fix two subgradients of $r$ at these points: The subgradient $n_{x_{\mathtt{OPT}}}\in\partial r(x_{\mathtt{OPT}})$ is chosen such that $f_{0}(x)+r(x_{\mathtt{OPT}})+\langle n_{x_{\mathtt{OPT}}},x-x_{\mathtt{OPT}}\rangle$ is minimized over $f_{s}(x)\leq 0$ at $x_{\mathtt{OPT}}$ for all $s=1\ldots m$. The subgradient $n_{x_{\mathtt{SL}}}\in\partial r(x_{\mathtt{SL}})$ can be chosen freely.
These two reference subgradients of $r$ facilitate considering two lower bounds of the objective $f_{0}+r$ for either $y\in\{x_{\mathtt{OPT}},x_{\mathtt{SL}}\}$, denoted by
@@
At each iteration $k$, we denote the relative difference between $x_{k}$ and $y\in\{x_{\mathtt{OPT}},x_{\mathtt{SL}}\}$ in (relaxed) objective value or feasibility on the selected constraint function $f_{s(x_{k})}$ by
@@
Note $\delta_{k}(y)$ is always finite since the real-valued objective function lower bound $h_{y}$ is used instead of $f_{0}+r$ which takes value in the extended reals. Indeed, consider $r$ as an indicator function for a simple constraint set. This set is projected onto at each iteration where $x_{k}$ is feasible for all of the functional constraints, ensuring $x_{k+1}$ is feasible for the simple constraint. We cannot, however, guarantee that $x_{k}$ satisfies the simple constraint, and so $r(x_{k})$ may be infinite.
The sign of $\delta_{k}(y)$ may vary. If $y=x_{\mathtt{OPT}}$, then $\delta_{k}(x_{\mathtt{OPT}})$ is nonnegative, being lower bounded by the level of suboptimality or current infeasibility
$\delta_{k}(x_{\mathtt{OPT}})\geq\begin{cases}h_{x_{\mathtt{OPT}}}(x_{k})-p_{\star}&\mathrm{if\ }x_{k}\mathrm{\ is\ feasible}\\ f_{s(x_{k})}(x_{k})&\mathrm{otherwise}\end{cases}\ \geq 0\ .$
When $y=x_{\mathtt{SL}}$ and $x_{k}$ is feasible, $\delta_{k}(y)$ may be negative but is bounded below by
$\delta_{k}(x_{\mathtt{SL}})\geq\inf h_{x_{\mathtt{SL}}}-h_{x_{\mathtt{SL}}}(x_{\mathtt{SL}})>-\infty\ .$
When $y=x_{\mathtt{SL}}$ and $x_{k}$ is infeasible, $\delta_{k}(y)$ is strictly positive, being bounded below by
$\delta_{k}(x_{\mathtt{SL}})\geq 0-\max_{s=1\dots m}f_{s}(x_{\mathtt{SL}})>0\ .$
@@
Assumption C. For both $y\in\{x_{\mathtt{OPT}},x_{\mathtt{SL}}\}$, there exist constants $L_{0}$, $L_{1}{\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}\geq 0}$ such that every iterate $x_{k}$ has
@@
This assumption captures several common settings. The following lemma shows this condition holds for any additive combination of nonsmooth Lipschitz and smooth settings with bounded variance in the stochastic subgradient oracles.
Lemma 2.4.
If there exists functions $f^{(1)}_{s},f^{(2)}_{s}$ such that each $f_{s}=f_{s}^{(1)}+f^{(2)}_{s}$ for $s=0,\dots,m$ where $f^{(1)}_{s}$ is uniformly $M$-Lipschitz and $f^{(2)}_{s}$ has uniformly $L$-Lipschitz gradient and $g_{s(x_{k})}(x_{k};\xi_{k})$ has variance uniformly bounded by $\sigma^{2}$, then Assumption C holds.
The proof of this lemma is deferred to the appendix where the explicit constants $L_{0}$ and $L_{1}$ can be found. This established that in the standard nonsmooth optimization setting where each $f_{s}$ is uniformly Lipschitz, Assumption C holds with $L_{1}=0$. Further, Assumption C also holds in the standard smooth optimization setting where each $f_{s}$ has uniformly $L$-Lipschitz gradient. Section 4 gives an illustrative numerical example of this form where Assumption C holds despite the objective being neither Lipschitz nor smooth.
@@
$T_{0}$ $:=\sup\left\{k\in\{0,1,2\dots\}\mid L_{1}\alpha_{k}>1\right\}\ ,$ (2.9)
$C_{0}$ $:=\sum_{k=0}^{T_{0}}\lambda_{k}\max\left\{L_{1}\alpha_{k}-1,0\right\}\max\{\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}}),\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{SL}})\}\ .$ (2.10)
Observe that $T_{0}$, and hence $C_{0}$, is bounded if $\alpha_{k}$ is eventually always less or equal to $1/L_{1}$, capturing all stepsize policies with $\alpha_{k}\rightarrow 0$. Despite being bounded, the constant $C_{0}$ can be exponentially large in $T_{0}$. The toy example considered above has $T_{0}=397$ and $C_{0}>10^{112}$. Such potentially exponential-sized constants can be avoided through careful stepsize selection as if one selects $\alpha_{k}\leq 1/L_{1}$ for all $k$ as then $T_{0}=-\infty$ and $C_{0}=0$. In particular, under the classic assumption that subgradients seen are uniformly bounded, Assumption C holds with $L_{1}=0$ and so $C_{0}=0$. Regardless, our convergence guarantees apply whenever $C_{0}$ is finite. As a natural consequence of our main convergence analysis, we find the rate $x_{k}$ can diverge, and hence the constant $C_{0}$, are at most exponential in $T_{0}$ (see Proposition 3.4).
@@ 3 Primal-Dual Equivalence and Convergence Analysis
Let $\{x_{k}\}$ be the sequence of iterates of the primal method (2.2) with stepsizes $\alpha_{k}>0$ and $\{y_{k}\}$ be the sequence of iterates of the dual method (2.3) for some $\lambda_{k},\beta_{k},\mu\geq 0$. If $x_{0}=y_{0}$, $\beta_{k}=\bar{\beta}\geq 0$ is constant, and
$\alpha_{k}=\frac{\lambda_{k}}{\mu\sum_{i=0}^{k}\lambda_{i}+\bar{\beta}}\ ,$ (3.1)
then these methods are equivalent, that is $x_{k}=y_{k}$.
@@
We prove this by inductively showing that $x_{k}=y_{k}$ and for any iteration with $s(x_{k})=0$ that $n_{k+1}=\frac{1}{\alpha_{k}}(x_{k}-\alpha_{k}g_{0}(x_{k};\xi_{k})-x_{k+1})$ (i.e., the subgradients of $r$ produced by the dual method are exactly those produced by the primal method via (2.1)). By assumption, $x_{0}=y_{0}$. Suppose for induction that $x_{i}=y_{i}$ for all $i=0,\dots,k$ and for all $i=0,\dots,k-1$ with $s(x_{i})=0$ that $n_{i+1}=\frac{1}{\alpha_{i}}(x_{i}-\alpha_{i}g_{0}(x_{i};\xi_{i})-x_{i+1})$. Let $\bar{g}_{i}=g_{0}(y_{i};\xi_{i})+n_{i+1}$ if $y_{i}$ is feasible and $g_{s(y_{i})}(y_{i};\xi_{i})$ otherwise. For each $i<k$, we claim $x_{i+1}=x_{i}-\alpha_{i}\bar{g}_{i}$. If $x_{i}$ is infeasible, this is immediate. If $x_{i}$ is feasible, rearrangement of our inductive hypothesis of on $n_{i+1}$ gives the claim. Inductively, we conclude for $i<k$ that $x_{i+1}-x_{0}=-\sum_{t=0}^{i}\alpha_{t}\bar{g}_{t}.$
By Lemma 2.1, $y_{k+1}$ is the unique solution to $\sum_{i=0}^{k}\lambda_{i}\left(\bar{g}_{i}+\mu(y_{k+1}-y_{i})\right)+\bar{\beta}(y_{k+1}-y_{0})=0.$ Rearranging and simplifying this, it follows that
$y_{k+1}$ $=y_{0}+\frac{\sum_{i=0}^{k}\lambda_{i}\mu(y_{i}-y_{0})-\sum_{i=0}^{k}\lambda_{i}\bar{g}_{i}}{\sum_{i=0}^{k}\lambda_{i}\mu+\bar{\beta}}$
@@
$=y_{0}-\frac{\sum_{t=0}^{k-1}\lambda_{t}\bar{g}_{t}\frac{\sum_{i=t+1}^{k}\mu\lambda_{i}}{\sum_{i=0}^{t}\mu\lambda_{{\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}i}}+\bar{\beta}}+\sum_{i=0}^{k}\lambda_{i}\bar{g}_{i}}{\sum_{i=0}^{k}\lambda_{i}\mu+\bar{\beta}}$
$=y_{0}-\frac{\sum_{t=0}^{k-1}\lambda_{t}\bar{g}_{t}\left(\frac{\sum_{i=t+1}^{k}\mu\lambda_{i}}{\sum_{i=0}^{t}\mu\lambda_{{\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}i}}+\bar{\beta}}+1\right)}{\sum_{i=0}^{k}\lambda_{i}\mu+\bar{\beta}}-\frac{\lambda_{k}\bar{g}_{k}}{\sum_{i=0}^{k}\lambda_{i}\mu+\bar{\beta}}$
$=y_{0}-\sum_{t=0}^{k-1}\frac{\lambda_{t}\bar{g}_{t}}{\sum_{i=0}^{t}\lambda_{{\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}i}}\mu+\bar{\beta}}-\frac{\lambda_{k}\bar{g}_{k}}{\sum_{i=0}^{k}\lambda_{i}\mu+\bar{\beta}}$
@@
where the second equality uses the inductive hypothesis for each $y_{i}-y_{0}=x_{i}-x_{0}=-\sum_{t=0}^{i-1}\alpha_{t}\bar{g}_{t}$, the third exchanges summands, and the remainder combines and simplifies terms. If $x_{k}$ is infeasible, it is immediate that $x_{k+1}=x_{k}-\alpha_{k}\bar{g}_{k}=y_{k+1}$. If $x_{k}$ is feasible, the above equality ensures
$\frac{1}{\alpha_{k}}(x_{k}-\alpha_{k}g_{0}(x_{k};\xi_{k})-y_{k+1})=n_{k+1}\in\partial r(y_{k+1})\ .$
Noting by (2.1), $x_{k+1}$ is the unique solution to $\frac{1}{\alpha_{k}}(x_{k}-\alpha_{k}g_{0}(x_{k};\xi_{k})-z)\in\partial r(z)$, we must have $x_{k+1}{\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}=x_{k}-\alpha_{k}(g_{k}+n_{k+1})}=y_{k+1}$ and $n_{k+1}=\frac{1}{\alpha_{k}}(x_{k}-\alpha_{k}g_{0}(x_{k};\xi_{k})-x_{k+1})$ as required. ∎
@@
Hence, provided dual weights $\lambda_{k}$, one can easily construct $\alpha_{k}$ as stated in the theorem, and conversely, given stepsizes $\alpha_{k}$, one can easily construct corresponding weights $\lambda_{k}$. For example, fixing $\lambda_{0}=1$, $\bar{\beta}=0$ multipliers for several common stepsizes are
@@ 3.1 Statement of Primal-Dual Convergence Guarantees
$\mathtt{primal\mbox{-}gap}_{T}:=\frac{\sum_{k<T:s(x_{k})=0}\lambda_{k}h_{x_{\mathtt{OPT}}}(x_{k})}{\sum_{k<T:s(x_{k})=0}\lambda_{k}}-p_{\star}\ ,$ (3.3)
@@
$\mathtt{dual\mbox{-}gap}_{T}:=p_{\star}-\inf\frac{M^{(T-1)}}{\sum_{k<T:s(x_{k})=0}\lambda_{k}}\ ,$ (3.4)
@@
Under Assumptions A-C, for any primal stepsizes $\alpha_{k}>0$ and dual weights $\lambda_{k}>0$ satisfying (3.1) with $\bar{\beta}=0$, the stochastic switching proximal subgradient method (2.2) has
$\mathbb{E}_{\xi}\left[\frac{\sum_{k<T:s(x_{k})=0}\lambda_{k}}{\sum_{k=0}^{T-1}\lambda_{k}}\right]\geq\frac{\tau_{\mathtt{SL}}}{2(h_{x_{\mathtt{SL}}}(x_{\mathtt{SL}})-\inf h_{x_{\mathtt{SL}}})+\tau_{\mathtt{SL}}}\left(1-\frac{L_{0}^{2}\sum_{k=0}^{T-1}\lambda_{k}\alpha_{k}+C_{0}}{{\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}\tau_{\mathtt{SL}}}\sum_{k=0}^{T-1}\lambda_{k}}\right)$
@@
Under Assumptions A-C, for any primal stepsizes $\alpha_{k}>0$ and dual weights $\lambda_{k}>0$ satisfying (3.1) with $\bar{\beta}=0$, the stochastic switching proximal subgradient method (2.2) has
@@
Theorem 3.2 recovers the primal convergence rate of [11, Section 3.2] with $\alpha_{k}=2/(\mu(k+2))$ and extends it to be a primal-dual guarantee covering proximal, switching, and non-Lipschitz settings. Theorem 3.1 shows this stepsize corresponds to dual averaging with weights $\lambda_{k}=k+1$ and $\bar{\beta}=0$. When $m=0$, Theorem 3.2 ensures
@@
using that $\sum_{k=0}^{T-1}\lambda_{k}=T(T+1)/2$ and $\sum_{k=0}^{T-1}\lambda_{k}\alpha_{k}=2(T-\sum_{k=2}^{T+1}1/k)/\mu\leq 2T/\mu$. When $m>0$ and the subgradient oracle is deterministic, applying Proposition 3.1 gives a rate worse by only a factor depending on the Slater point of
$\mathtt{primal\mbox{-}gap}_{T}+\mathtt{dual\mbox{-}gap}_{T}+\frac{\mu}{2}\|x_{T}-x_{\mathtt{OPT}}\|_{2}^{2}$
$\qquad\qquad\leq\frac{2(h_{x_{\mathtt{SL}}}(x_{\mathtt{SL}})-\inf h_{x_{\mathtt{SL}}})+\tau_{\mathtt{SL}}}{{\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}\tau_{\mathtt{SL}}-\frac{4L_{0}^{2}}{\mu(T+1)}-\frac{2C_{0}}{T(T+1)}}}\left(\frac{4L_{0}^{2}}{\mu(T+1)}+\frac{2C_{0}}{T(T+1)}\right)\ .$
Here, the role of $C_{0}$, defined in (2.10), bounding the effect of the non-Lipschitz constant $L_{1}$ becomes clear. The only role $L_{1}$ plays in our rate is via $T_{0}$, which in turn $C_{0}$ may be exponentially large in (see Proposition 3.4). If $C_{0}$ is small, then convergence will be dominated by the classic $O(L_{0}^{2}/\mu T)$ term as the dependence on $C_{0}$ shrinks at a fast $O(1/T^{2})$ rate.